# Rectified Diffusion Guidance for Conditional Generation

### v1

@@ A.2 Proof of Theorem 2
$\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c)$ $=\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ (S41)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=-\frac{1}{2}\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c)$ (S43)
@@
Since $02k-\gamma<$ 0 for $k=0,1,\cdots,n$, we have $(T+1)^{\frac{2k-\gamma}{2}}\rightarrow 0$ as $T$ goes to infinity and hence
@@ A.3 Proof of Theorem 3
Similarly, by $p_{\theta,\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t}|c)=\int q_{0}(\mathbf{x}_{0}|c)q_{0T}(\mathbf{x}_{T}|\mathbf{x}_{0},c)p_{\theta,\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t}|\mathbf{x}_{T},c)\mathrm{d}\mathbf{x}_{0}\mathrm{d}\mathbf{x}_{T}$ we have
@@
$=\frac{\sigma_{t-1}}{\sigma_{t}}\Delta_{t}-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},t)]).$ (S110)
$\Delta_{t}=0$ implies that $\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]=\mathbb{E}_{p_{\theta,\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t}|c)}[\tilde{\mathbf{x}}_{t}]$. Therefore, by Eqs. S107 and S108 we have $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]=\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},c,t)]$. According to Lemma 2 and by calculating the expectation over $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{t}$ respectively, we have
@@
$\Delta_{t-1}=-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\mathbf{x}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]).$ (S113)
@@ A.4 Proof of Theorem 4
$\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)$ $=\gamma_{1}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+\gamma_{0}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ (S114)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=-\frac{1}{2}\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)$ (S116)
@@ A.5 Proof of Theorem 5
According to Eqs. 29 and 30, we can write the variational lower bound of $p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0:T}|c)$ as below:
$J_{\delta,\gamma_{1},\gamma_{0}}$ $=\mathbb{E}_{q_{\delta}(\mathbf{x}_{0:T}|c)}[\log q_{\delta}(\mathbf{x}_{1:T}|\mathbf{x}_{0},c)-\log p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0:T}|c)]$ (S123)
$=\mathbb{E}\left[-\log p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|\mathbf{x}_{1},c)\right]$
$\qquad+\mathbb{E}\left[\sum_{t=2}^{T}D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))\right]$
@@
$\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))]$ (S125)
$=$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{t},c),c))]$ (S126)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{t},c)\|_{2}^{2}]$ (S127)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[-\log p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|\mathbf{x}_{1},c))]$ (S135)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{1},c)\|_{2}^{2}]+C_{3}$ (S136)

### v2

@@ A.2 Proof of Theorem 2
$s_{t,\gamma}(\mathbf{x}_{t},c)$ $=\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ (S41)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=-\frac{1}{2}s_{t,\gamma}(\mathbf{x}_{t},c)$ (S43)
@@
Since $2k-\gamma<0$ for $k=0,1,\cdots,n$, we have $(T+1)^{\frac{2k-\gamma}{2}}\rightarrow 0$ as $T$ goes to infinity and hence
@@ A.3 Proof of Theorem 3
Similarly, by $\hat{p}_{\theta}(\tilde{\mathbf{x}}_{t}|c)=\int q_{0}(\mathbf{x}_{0}|c)q_{0T}(\mathbf{x}_{T}|\mathbf{x}_{0},c)\hat{p}_{\theta}(\tilde{\mathbf{x}}_{t}|\mathbf{x}_{T},c)\mathrm{d}\mathbf{x}_{0}\mathrm{d}\mathbf{x}_{T}$ we have
@@
$=\frac{\sigma_{t-1}}{\sigma_{t}}\Delta_{t}-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},t)])$ (S110)
$=\frac{\sigma_{t-1}}{\sigma_{t}}\Delta_{t}-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[\bm{\epsilon}_{\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t})]).$ (S111)
$\Delta_{t}=0$ implies that $\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]=\mathbb{E}_{\hat{p}_{\theta}(\tilde{\mathbf{x}}_{t}|c)}[\tilde{\mathbf{x}}_{t}]$. Therefore, by Eqs. S107 and S108 we have $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]=\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},c,t)]$. According to Lemma 2 and by calculating the expectation over $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{t}$ respectively, we have
@@
$\Delta_{t-1}$ $=-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\mathbf{x}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)])$ (S114)
$=-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\gamma_{1},\gamma_{0}}(\mathbf{x}_{t})].$ (S115)
@@ A.4 Proof of Theorem 4
$s_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t},c)$ $=\gamma_{1}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+\gamma_{0}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ (S116)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=-\frac{1}{2}s_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t},c)$ (S118)
@@ A.5 Proof of Theorem 5
According to Eqs. 32 and $3$, we can write the variational lower bound of $\hat{p}_{\theta}(\mathbf{x}_{0:T}|c)$ as below:
$J_{\delta,\gamma_{1},\gamma_{0}}$ $=\mathbb{E}_{q_{\delta}(\mathbf{x}_{0:T}|c)}[\log q_{\delta}(\mathbf{x}_{1:T}|\mathbf{x}_{0},c)-\log\hat{p}_{\theta}(\mathbf{x}_{0:T}|c)]$ (S125)
$=\mathbb{E}\left[-\log\hat{p}_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},c)\right]$
$\qquad+\mathbb{E}\left[\sum_{t=2}^{T}D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))\right]$
@@
$\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))]$ (S127)
$=$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c),c))]$ (S128)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c)\|_{2}^{2}]$ (S129)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[-\log\hat{p}_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},c))]$ (S137)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{1},c)\|_{2}^{2}]+C_{3}$ (S138)
@@ 10 Returns:
36 # Record the mean coefficient for other conditions.
37 coeffs.update(avg=sum(coeffs.values()) / len(coeffs)
38